# Model Merging in the Era of Large Language Models: Methods, Applications, and Future Directions

### 10 Mar 2026 v1

@@ 6 Application Scenarios and Practical Impact
We now shift focus from how models are merged to where and why. This section addresses the Scenarios dimension of our FUSE taxonomy, organizing applications by their underlying motivation. We examine three categories: capability-driven applications that leverage merging for enhanced performance and generalization, alignment-driven applications that target safety and value conformance, and efficiency-driven applications motivated by computational, communication, or resource constraints in settings such as federated learning and model compression.
@@ 6.1 Capability Augmentation Through Multi-Task and Multilingual Merging
Model merging provides a practical pathway for capability augmentation: rather than training monolithic models on aggregated datasets or maintaining costly ensembles, practitioners can compose specialized abilities by merging independently fine-tuned models into a unified system (Ilharco et al., 2023).
@@
The theoretical foundation for multi-task capability merging rests on the observation that fine-tuned models sharing a common pretrained initialization occupy proximate regions in weight space, enabling meaningful interpolation between their specialized parameters. When multiple models are fine-tuned for distinct tasks, such as summarization, question answering, and code generation, the resulting task vectors encode complementary capabilities that can be arithmetically combined. The merged model $\theta_{\text{merged}}=\theta_{\text{base}}+\sum_{i=1}^{n}\lambda_{i}\tau_{i}$, where $\tau_{i}=\theta_{\text{fine-tuned},i}-\theta_{\text{base}}$ represents the task vector for task $i$ and $\lambda_{i}$ denotes its scaling coefficient, ideally inherits the union of capabilities from all constituent models. This formulation, introduced by Ilharco et al. (2023), demonstrated that task vector addition yields models capable of performing multiple tasks simultaneously, with experiments on LLMs showing that merging multiple task vectors maintained an average accuracy within 2-3% of individually fine-tuned models. Subsequent work (Yadav et al., 2023; Yu et al., 2023) demonstrated that carefully guided task vector merging achieves strong performance across natural language processing tasks, validating the practical viability of this mathematical formulation.
@@
However, naïve task vector addition often suffers from parameter interference when merging multiple models, where redundant or conflicting parameter updates across tasks degrade performance. Several sophisticated methods address this challenge. TIES-Merging (Yadav et al., 2023) introduces a three-step procedure, trimming low-magnitude parameters, electing sign consensus, and merging only aligned parameters, which reduces interference by trimming, sign election, and aligned merging on T5 (Raffel et al., 2020)-based multi-task benchmarks. DARE (Yu et al., 2023) takes a complementary approach by randomly dropping delta parameters with high probability and rescaling remaining values, exploiting the observation that fine-tuning induces sparse, redundant changes; this method enabled successful merging of up to six specialized LLMs while preserving over 90% of individual task performance. More recent approaches such as Model Stock (Jang et al., 2024) leverage geometric insights about the fine-tuning loss landscape, demonstrating that averaging models from the vicinity of the optimal weight region yields superior merged models with only two to three fine-tuned constituents required. EMR-Merging (Huang et al., 2024) achieves tuning-free high-performance model merging through elastic mask re-weighting. Twin-Merging (Lu et al., 2024) introduces dynamic integration of modular expertise, while PCB-Merging (Du et al., 2024) addresses parameter competition (Du et al., 2024) balancing. AdaMerging (Yang et al., 2024d) provides adaptive model merging specifically designed for multi-task learning (Aghajanyan et al., 2021) scenarios.
@@
Instruction tuning consolidation represents a particularly impactful application of capability merging in the large language model ecosystem. Individual instruction-tuned models such as WizardLM (Xu et al., 2024; Yang et al., 2024d) often excel on specific instruction formats or task categories while exhibiting degraded performance on others. By merging multiple instruction-tuned variants, each optimized for different instruction styles, complexity levels, or domain coverage, practitioners can construct unified models that demonstrate robust instruction-following across diverse prompting scenarios. Empirical studies demonstrate the efficacy of this approach: reported methods have achieved 99.4% of the performance of traditional multi-task training on instruction-following benchmarks while requiring no additional training. Similarly, experiments merging WizardLM, WizardMath (Luo et al., 2023), and code-specialized variants such as Code Llama (Rozière et al., 2023) showed 8-12% improvements on aggregate benchmarks compared to individual constituents (Li et al., 2023a). LM-Cocktail (Xiao et al., 2023) further demonstrates resilient tuning through strategic model merging. FuseLLM (Wan et al., 2024a) takes a complementary approach by fusing the knowledge of multiple structurally diverse LLMs through lightweight continual training on probability distributions generated by the source models, enabling cross-architecture knowledge transfer that traditional weight-space methods cannot achieve. This consolidation approach is widely used in community-driven model development, where independent contributors fine-tune base models on complementary instruction datasets.
@@
Multilingual capability transfer through merging addresses the challenge of building language models that perform well across typologically diverse languages. Conventional approaches require either massive multilingual pretraining corpora or accept degraded performance on low-resource languages. Model merging offers an alternative pathway: specialized models fine-tuned on individual languages or language families can be combined to transfer cross-lingual capabilities (Huang et al., 2023b). Huang et al. (2023b) demonstrated that identifying and preserving language-specific neuronal subnetworks during merging enables effective cross-lingual transfer without catastrophic forgetting of source language capabilities. Chat Vector (Huang et al., 2023b; b) provides a simple yet effective approach to equip LLMs with instruction-following and model alignment capabilities in new languages through task vector manipulation. Complementary work on cross-lingual model merging (Huang et al., 2023b) showed that interpolating between English-specialized and target-language models improved low-resource language performance by 15-20% on XNLI benchmarks compared to English-only fine-tuned baselines. This proves particularly effective when merging models that share vocabulary coverage but differ in their language-specific fine-tuning. Recent investigations into language-agnostic and language-specific parameter separation suggest that modular approaches, merging language-specific adapters rather than full models, offer finer-grained control over cross-lingual capability composition while mitigating interference between typologically distant language pairs.
@@
The practical impact of capability merging is evidenced by its adoption in open-source communities, where merged models have demonstrated measurable performance gains on standardized evaluation suites. For instance, analysis of Open LLM Leaderboard submissions revealed that merged models constituted a substantial portion of top-performing entries, with specific combinations such as instruction-tuned and reasoning-specialized merges achieving 3-7% absolute improvements over their strongest constituents (Yadav et al., 2024a). However, capability merging exhibits well-documented failure modes that practitioners must consider. Negative transfer occurs most frequently when constituent tasks impose conflicting optimization pressures, for example, merging models fine-tuned for concise summarization with those trained for detailed elaboration often yields degraded performance on both objectives (Yadav et al., 2023). Empirical analyses (Yadav et al., 2023) identified that task dissimilarity, measured through gradient conflict metrics, strongly predicts merge quality degradation, with highly dissimilar task pairs showing up to 25% performance drops compared to individual models. Furthermore, (Li et al., 2023a) demonstrated that representation bias toward dominant tasks emerges when merging models with imbalanced training scales, requiring explicit representation alignment to mitigate. These findings underscore that successful capability augmentation depends critically on careful task selection and appropriate interference mitigation strategies. While such capability enhancement represents the predominant application of model merging, task vector arithmetic and parameter combination extend naturally to improving alignment with human values and mitigating safety risks.
@@
AdapterSoup (Chronopoulou et al., 2023) demonstrates that averaging domain-specific adapters can surpass any individual adapter when the target domain overlaps with multiple training domains, outperforming individual adapter selection on held-out domains. LoRAHub (Huang et al., 2023a) extends this principle to few-shot scenarios, composing LoRA modules through gradient-free Bayesian optimization over combination weights. Critically, these adapter-level approaches reduce the merge cost from $\mathcal{O}(|\theta|)$ to $\mathcal{O}(r\cdot d)$ where $r\ll d$ is the adapter rank, making iterative composition practical even for 70B-parameter models.
@@ 6.2 Alignment, Safety, and Bias Mitigation Applications
Model merging has also emerged as a cost-effective paradigm for improving alignment with human values and mitigating safety risks in deployed language models. The fundamental insight motivating these approaches is that preference-aligned models, typically trained through reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022), or direct preference optimization (DPO) (Rafailov et al., 2023), can be combined with capable base models to achieve favorable tradeoffs between helpfulness (Liang et al., 2025) and harmlessness without incurring the substantial computational costs of full alignment training. This approach complements constitutional methods by enabling post-hoc alignment adjustment through parameter-space operations rather than requiring extensive retraining with curated feedback.
@@
The task vector formalism provides a particularly elegant mechanism for targeted bias mitigation (Ilharco et al., 2023). Given a base model with parameters $\theta_{\text{base}}$ and a model fine-tuned to exhibit specific biased behaviors $\theta_{\text{biased}}$, the bias task vector $\tau_{\text{bias}}=\theta_{\text{biased}}-\theta_{\text{base}}$ encodes the parametric representation of the undesired behavior. Through task vector negation, a debiased model can be constructed as $\theta_{\text{debiased}}=\theta_{\text{base}}-\alpha\cdot\tau_{\text{bias}}$, where the scaling coefficient $\alpha$ controls the intensity of bias removal. This subtractive approach enables surgical removal of specific behavioral patterns, including toxic language generation, demographic stereotyping, and harmful content production, without degrading the model’s core linguistic competencies. However, a significant practical limitation of this formulation is the requirement for an explicitly fine-tuned $\theta_{\text{biased}}$ model that isolates the undesired behavior. Recent work addresses this challenge through probe-based bias identification and activation steering methods, which enable bias vector estimation without requiring explicit fine-tuning on harmful objectives (Li et al., 2023a).
@@
Beyond task arithmetic, alternative merging methodologies offer distinct mechanisms for alignment preservation. TIES-Merging addresses parameter interference through sign-based conflict resolution, which proves particularly valuable when merging models aligned on potentially conflicting value dimensions by preserving the dominant alignment signal in each parameter (Yadav et al., 2023). Fisher-weighted merging leverages parameter importance estimates to prioritize safety-critical weights during combination, enabling assign higher Fisher information scores to parameters associated with alignment behaviors (Li et al., 2023a). DARE employs random parameter pruning before merging, which empirical studies suggest can reduce the propagation of misaligned behaviors embedded in redundant parameters (Yu et al., 2023).
@@
The merging of RLHF-trained models presents unique opportunities and challenges for alignment enhancement. When multiple models have been independently aligned using different preference datasets or reward models, their combination can potentially yield more robust alignment that generalizes across diverse value dimensions (Ram’e et al., 2024). The interpolation $\theta_{\text{merged}}=(1-\lambda)\theta_{\text{helpful}}+\lambda\theta_{\text{harmless}}$ between helpfulness-optimized and harmlessness-optimized variants enables practitioners to navigate the alignment tradeoff frontier (Ram’e et al., 2024; 2023b), selecting operating points appropriate for specific deployment contexts. Empirical investigations reveal that intermediate interpolation coefficients often outperform both endpoints on aggregate preference metrics, suggesting that merged models capture complementary aspects of human preferences that neither constituent model fully represents (Ram’e et al., 2024). This finding has been corroborated across multiple model families (Li et al., 2023a).
@@
Evaluating alignment preservation in merged models requires multi-faceted assessment protocols spanning automated metrics, benchmark evaluations, and adversarial testing. Standard safety benchmarks provide quantitative measures of factuality, toxicity, and social bias respectively. Red-teaming protocols systematically probe merged models for adversarial vulnerabilities that may emerge from parameter combination, while automated red-teaming approaches enable scalable safety evaluation (Wan et al., 2024b). Reward model scoring on held-out preference data offers another evaluation dimension, with recent work proposing specialized metrics for detecting alignment degradation after merging. Notably, evaluation must assess not only aggregate safety metrics but also consistency of alignment across capability levels, as merged models may exhibit capability-dependent safety profiles (Djuhera et al., 2025).
@@
Despite these advances, alignment-focused merging faces several known limitations that warrant careful consideration. First, the capability-safety tradeoff proves difficult to navigate in practice: aggressive bias removal via task vector negation frequently degrades model fluency and task performance, while conservative merging may inadequately address safety concerns (Ram’e et al., 2024). Second, isolating clean bias vectors remains challenging when biased behaviors are entangled with legitimate capabilities, for instance, removing gender bias associations may inadvertently impair performance on gender-related factual queries. Third, validation presents fundamental difficulties: merged models may exhibit emergent unsafe behaviors not present in constituent models, yet comprehensive safety evaluation across all possible inputs remains intractable. Finally, the alignment properties of merged models may be more brittle than those of directly trained models, showing greater susceptibility to jailbreaking attacks (Djuhera et al., 2025).
@@
Safety-aware merging strategies have gained prominence in addressing adversarial vulnerabilities of merged models. Recent work on mitigating backdoor effects in multi-task model merging through safety-aware subspace identification (Djuhera et al., 2025) highlights growing recognition that naive merging may inadvertently propagate or amplify security vulnerabilities present in constituent models. SafeMERGE (Djuhera et al., 2025) proposes selective layer-wise model merging to preserve safety alignment in fine-tuned LLMs. These approaches project merged parameters onto subspaces that preserve safety-critical behaviors while enabling capability transfer, often leveraging activation analysis to identify safety-relevant parameter subspaces (Djuhera et al., 2025). WARM (Ram’e et al., 2024) demonstrates the benefits of weight averaged reward models for improving alignment robustness. This represents a principled integration of security considerations into the merging pipeline.
The healthcare domain exemplifies the critical importance of alignment-preserving merging strategies, where domain-specialized models must maintain stringent safety properties including factual accuracy, appropriate uncertainty quantification, and avoidance of harmful medical advice (Li et al., 2023a). To address these requirements, researchers have developed uncertainty-aware merging protocols that combine clinical fine-tuned models with safety-aligned base models using Fisher-weighted averaging, prioritizing parameters associated with calibrated uncertainty estimation. Additionally, factual consistency constraints during merging, implemented through knowledge-preserving regularization terms, help ensure that clinical expertise transfer does not compromise factual guardrails established during alignment training. Domain-specific evaluation protocols for merged medical models now incorporate clinical accuracy benchmarks alongside standard safety metrics, enabling comprehensive assessment of the capability-safety balance (Li et al., 2023a).
@@
Similarly, cybersecurity applications demand rigorous attention to the security implications of model merging, as merged models may exhibit unexpected vulnerabilities arising from the combination of individually benign model components (Li et al., 2023a). This observation has spurred investigation into adversarial robustness preservation during merging, with emerging techniques incorporating explicit security constraints into the merging objective. Specifically, robust merging protocols for security-critical applications employ adversarial example validation during coefficient tuning, rejecting merge configurations that degrade performance on curated attack datasets. Furthermore, differential testing between merged and constituent models enables detection of emergent vulnerabilities unique to the merged configuration (Li et al., 2023a). These domain-specific considerations become even more complex when merging must occur across organizationally or geographically distributed settings, where data privacy constraints and communication limitations introduce additional challenges that significantly reshape how model combination can be performed.
@@
The same arithmetic that enables safety injection also enables safety removal: subtracting the alignment task vector $\tau_{\text{safe}}$ from a safety-tuned model can recover an unaligned variant, raising concerns about misuse (Ilharco et al., 2023). Furthermore, reward model merging via WARM (Ram’e et al., 2024) has shown that averaging independently trained reward models reduces reward hacking susceptibility by 4–7% on the RewardBench suite, but the optimal number of reward models to merge and the conditions under which averaging preserves reward calibration remain theoretically uncharacterized. These dual-use considerations necessitate community-level governance frameworks for responsible model merging.

### 30 Mar 2026 v2

@@ 6 Application Scenarios and Practical Impact
We now shift focus from how models are merged to where and why. This section covers the Scenarios dimension of our FUSE taxonomy, organizing applications by their underlying motivation. We examine three categories. These include capability-driven applications that use merging for enhanced performance and generalization, alignment-driven applications that target safety and value conformance, and efficiency-driven applications motivated by computational, communication, or resource constraints in settings such as federated learning and model compression.
@@ 6.1 Capability Augmentation Through Multi-Task and Multilingual Merging
Model merging provides a practical pathway for capability augmentation. Rather than training monolithic models on aggregated datasets or maintaining costly ensembles, practitioners can compose specialized abilities by merging independently fine-tuned models into a unified system (Ilharco et al., 2023).
@@
Multi-task capability merging rests on the theoretical observation that fine-tuned models sharing a common pretrained initialization occupy proximate regions in weight space, so that meaningful interpolation between their specialized parameters is possible. When multiple models are fine-tuned for distinct tasks, such as summarization, question answering, and code generation, the resulting task vectors encode complementary capabilities that can be arithmetically combined. The merged model $\theta_{\text{merged}}=\theta_{\text{pre}}+\sum_{i=1}^{n}\lambda_{i}\tau_{i}$, where $\tau_{i}=\theta_{i}-\theta_{\text{pre}}$ represents the task vector for task $i$ and $\lambda_{i}$ denotes its scaling coefficient, ideally inherits the union of capabilities from all constituent models. This formulation, introduced by Ilharco et al. (2023), showed that task vector addition yields models capable of performing multiple tasks simultaneously while maintaining strong performance on individual tasks. Subsequent work (Yadav et al., 2023; Yu et al., 2024) showed that carefully guided task vector merging produces strong performance across natural language processing tasks, validating the practical viability of this mathematical formulation.
@@
However, naïve task vector addition often suffers from parameter interference when merging multiple models, where redundant or conflicting parameter updates across tasks degrade performance. Several dedicated methods tackle this challenge. TIES-Merging (Yadav et al., 2023) introduces a three-step procedure that trims low-magnitude parameters, elects sign consensus via majority voting, and merges only aligned parameters, effectively reducing parameter interference across tasks. DARE (Yu et al., 2024) takes a complementary approach by randomly dropping delta parameters with high probability and rescaling remaining values, exploiting the observation that fine-tuning induces sparse, redundant changes; this method enabled successful merging of multiple specialized LLMs while largely preserving individual task performance. More recently, Model Stock (Jang et al., 2024b) uses geometric insights about the fine-tuning weight space, showing that layer-wise averaging of as few as two fine-tuned models can surpass methods such as Model Soups that require many more constituents. EMR-Merging (Huang et al., 2024a) attains tuning-free high-performance model merging through elect, mask, and rescale operations. Twin-Merging (Lu et al., 2024) introduces dynamic integration of modular expertise, while PCB-Merging (Du et al., 2024a) handles parameter competition balancing. AdaMerging (Yang et al., 2024d) provides adaptive model merging specifically designed for multi-task learning scenarios.
@@
Instruction tuning consolidation represents a particularly impactful application of capability merging in the large language model ecosystem. Individual instruction-tuned models such as WizardLM (Xu et al., 2024) often excel on specific instruction formats or task categories while exhibiting degraded performance on others. Disperse-Then-Merge (Fu et al., 2024) shows that dispersing instruction data across specialized models before merging pushes the limits of instruction tuning via alignment tax reduction. PAFT (Pentyala et al., 2024) introduces a parallel training framework that independently performs supervised fine-tuning and preference alignment (e.g., DPO) from the same pretrained model, then merges the resulting models through parameter fusing with delta sparsification to mitigate the alignment tax. By merging multiple instruction-tuned variants, each optimized for different instruction styles, complexity levels, or domain coverage, practitioners can construct unified models that exhibit robust instruction-following across diverse prompting scenarios. Empirical studies confirm the efficacy of this approach, with merged models posting performance competitive with traditional multi-task training on instruction-following benchmarks while requiring no additional training. Souper-Model (Maiti et al., 2025) shows that simple arithmetic operations on model weights can unlock strong LLM performance, validating the practical promise of merging. Similarly, community practitioners have reported that merging instruction-tuned variants such as WizardLM (Xu et al., 2024), WizardMath (Luo et al., 2025), and code-specialized models such as Code Llama (Rozière et al., 2023) can yield improvements on aggregate benchmarks compared to individual constituents (Yadav et al., 2025a). Zhang et al. (2024) propose unconstrained model merging that relaxes traditional constraints for enhanced LLM reasoning capabilities. ReasonAny (Yang et al., 2026a) further shows that reasoning capabilities can be incorporated into any model via simple and effective model merging. LM-Cocktail (Xiao et al., 2024) further shows resilient tuning through strategic model merging. FuseLLM (Wan et al., 2024) takes a complementary approach by fusing the knowledge of multiple structurally diverse LLMs through lightweight continual training on probability distributions generated by the source models, enabling cross-architecture knowledge transfer that traditional weight-space methods cannot achieve. FuseChat-3.0 (Yang et al., 2025c) extends this framework by incorporating preference optimization for heterogeneous chat model fusion. This consolidation approach is widely used in community-driven model development, where independent contributors fine-tune base models on complementary instruction datasets.
@@
Multilingual capability transfer through merging confronts the challenge of building language models that perform well across typologically diverse languages. Conventional approaches require either massive multilingual pretraining corpora or accept degraded performance on low-resource languages. Model merging offers an alternative pathway whereby specialized models fine-tuned on individual languages or language families can be combined to transfer cross-lingual capabilities. Chat Vector (Huang et al., 2024b) provides a simple yet effective approach to equip LLMs with instruction-following and model alignment capabilities in new languages through task vector manipulation, demonstrating that subtracting an English base model from its corresponding English chat model and adding the resulting vector to a model that has been continually pretrained on the target language effectively transfers conversational abilities without further training. Complementary work on cross-lingual model merging has shown that interpolating between English-specialized and target-language models can improve low-resource language performance markedly compared to English-only fine-tuned baselines (Bandarkar et al., 2025; Alexandrov et al., 2024). This proves particularly effective when merging models that share vocabulary coverage but differ in their language-specific fine-tuning. Recent investigations into language-agnostic and language-specific parameter separation suggest that modular approaches, merging language-specific adapters rather than full models, offer finer-grained control over cross-lingual capability composition while mitigating interference between typologically distant language pairs.
Multimodal model merging has gained momentum as researchers extend these techniques beyond text. Sung et al. (2023) provide a systematic empirical study identifying key success factors for multimodal merging, including initialization strategies and architectural choices. Model Composition for Multimodal Large Language Models (Chen et al., 2024a) combines vision and language capabilities through merging, while Model Tailor (Zhu et al., 2024a) combats catastrophic forgetting in multi-modal large language models through targeted merging strategies. Composing Parameter-Efficient Modules (Zhang et al., 2023) shows that arithmetic operations on LoRA parameters support flexible capability composition for LLMs. In open-source communities, merged models have accounted for a notable portion of top-performing Open LLM Leaderboard submissions, with instruction-tuned and reasoning-specialized merges achieving consistent improvements over their strongest constituents (Yadav et al., 2025a).
@@
Despite these successes, capability merging exhibits well-documented failure modes. Negative transfer occurs most frequently when constituent tasks impose conflicting optimization pressures. For example, merging models fine-tuned for concise summarization with those trained for detailed elaboration often degrades performance on both objectives (Yadav et al., 2023). Aakanksha et al. (2024) investigate whether data mixing or model merging is preferable, finding that the optimal strategy depends strongly on task similarity and data distribution characteristics. Empirical analyses (Yadav et al., 2023) confirm that gradient conflict metrics strongly predict merge quality degradation, with highly dissimilar task pairs showing marked performance drops. These findings highlight that successful capability augmentation depends on careful task selection and appropriate interference mitigation strategies. While such capability enhancement represents the predominant application of model merging, task vector arithmetic and parameter combination extend naturally to improving alignment with human values and mitigating safety risks.
@@
AdapterSoup (Chronopoulou et al., 2023) shows that averaging domain-specific adapters can surpass any individual adapter when the target domain overlaps with multiple training domains, outperforming individual adapter selection on held-out domains. LoRAHub (Huang et al., 2023) extends this principle to few-shot scenarios, composing LoRA modules through gradient-free optimization over combination weights. MerA (He et al., 2023) proposes merging pretrained adapters specifically for few-shot learning, showing that adapter-level merging supports effective task transfer with minimal data. Ostapenko et al. (2024) develop a framework for building and reusing a library of LoRAs toward modular LLMs, where adapters can be dynamically composed for novel tasks. LoRI (Zhang et al., 2025c) tackles the challenge of cross-task interference in multi-task low-rank adaptation by learning interference-resistant adapter representations. Critically, these adapter-level approaches reduce the merge cost from $\mathcal{O}(|\theta|)$ to $\mathcal{O}(r\cdot d)$ where $r\ll d$ is the adapter rank, making iterative composition practical even for 70B-parameter models.
@@ 6.2 Alignment, Safety, and Bias Mitigation Applications
Model merging also offers a cost-effective approach for improving alignment with human values and mitigating safety risks in deployed language models. The key insight motivating these approaches is that preference-aligned models, typically trained through reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022), or direct preference optimization (DPO) (Rafailov et al., 2023), can be combined with capable base models to achieve favorable tradeoffs between helpfulness and harmlessness (Yang et al., 2025a) without incurring the large computational costs of full alignment training. This approach complements constitutional methods by allowing post-hoc alignment adjustment through parameter-space operations rather than requiring extensive retraining with curated feedback.
@@
The task vector formalism provides a particularly elegant mechanism for targeted bias mitigation (Ilharco et al., 2023). Given a base model with parameters $\theta_{\text{pre}}$ and a model fine-tuned to exhibit specific biased behaviors $\theta_{\text{biased}}$, the bias task vector $\tau_{\text{bias}}=\theta_{\text{biased}}-\theta_{\text{pre}}$ encodes the parametric representation of the undesired behavior. Through task vector negation, a debiased model can be constructed as $\theta_{\text{debiased}}=\theta_{\text{pre}}-\alpha\cdot\tau_{\text{bias}}$, where the scaling coefficient $\alpha$ controls the intensity of bias removal. This subtractive approach enables surgical removal of specific behavioral patterns, including toxic language generation, demographic stereotyping, and harmful content production, without degrading the model’s core linguistic competencies. However, a notable practical limitation of this formulation is the requirement for an explicitly fine-tuned $\theta_{\text{biased}}$ model that isolates the undesired behavior. Alternative approaches based on probe-based bias identification and activation steering can enable bias vector estimation without requiring explicit fine-tuning on harmful objectives.
@@
Beyond task arithmetic, alternative merging methodologies offer distinct mechanisms for alignment preservation. TIES-Merging mitigates parameter interference through sign-based conflict resolution, which proves particularly valuable when merging models aligned on potentially conflicting value dimensions by preserving the dominant alignment signal in each parameter (Yadav et al., 2023). Fisher-weighted merging uses parameter importance estimates to prioritize safety-critical weights during combination, assigning higher Fisher information scores to parameters associated with alignment behaviors (Matena and Raffel, 2022). DARE applies random parameter pruning before merging, which empirical studies suggest can reduce the propagation of misaligned behaviors embedded in redundant parameters (Yu et al., 2024). Machine unlearning approaches (Liu et al., 2024b) show that merging-based techniques can accomplish targeted removal of unsafe knowledge from LLMs, while safety re-alignment methods (Bhardwaj et al., 2024) show that fine-tuning-induced safety degradation can be reversed through strategic parameter merging. LED-Merging (Ma et al., 2025) specifically targets safety-utility conflicts in model merging through gradient-based neuron locating, importance-based election, and conflicting update disjunction, allowing practitioners to navigate the tension between maintaining safety properties and enhancing task capabilities. Model Extrapolation Expedites Alignment (Zheng et al., 2025a) shows that extrapolating beyond interpolation in the direction of alignment improvements can accelerate convergence to desired safety properties.
@@
The merging of RLHF-trained models presents unique opportunities and challenges for alignment enhancement. When multiple models have been independently aligned using different preference datasets or reward models, their combination can potentially yield more robust alignment that generalizes across diverse value dimensions (Ram’e et al., 2024). The interpolation $\theta_{\text{merged}}=(1-\lambda)\theta_{\text{helpful}}+\lambda\theta_{\text{harmless}}$ between helpfulness-optimized and harmlessness-optimized variants enables practitioners to navigate the alignment tradeoff frontier (Ram’e et al., 2024; 2023b), selecting operating points appropriate for specific deployment contexts. Empirical investigations reveal that intermediate interpolation coefficients often outperform both endpoints on aggregate preference metrics, suggesting that merged models capture complementary aspects of human preferences that neither constituent model fully represents. Hu et al. (2025a) formalize this observation by showing that weight interpolation between pretrained and instruction-tuned models can simultaneously improve both alignment and calibration, reaching a Pareto-superior frontier that mitigates the alignment tax typically incurred by instruction tuning.
@@
Evaluating alignment preservation in merged models requires multi-faceted assessment protocols spanning automated metrics, benchmark evaluations, and adversarial testing. Standard safety benchmarks provide quantitative measures of factuality, toxicity, and social bias respectively. Red-teaming protocols systematically probe merged models for adversarial vulnerabilities that may emerge from parameter combination, while automated red-teaming approaches enable scalable safety evaluation. Reward model scoring on held-out preference data offers another evaluation dimension. In particular, evaluation must assess not only aggregate safety metrics but also consistency of alignment across capability levels, as merged models may exhibit capability-dependent safety profiles (Djuhera et al., 2025).
@@
Despite these advances, alignment-focused merging faces several known limitations that warrant careful consideration. First, the capability-safety tradeoff proves difficult to navigate in practice. Aggressive bias removal via task vector negation frequently degrades model fluency and task performance, while conservative merging may inadequately mitigate safety concerns (Ilharco et al., 2023). Second, isolating clean bias vectors remains challenging when biased behaviors are entangled with legitimate capabilities, for instance, removing gender bias associations may inadvertently impair performance on gender-related factual queries. Third, validation presents considerable difficulties, as merged models may exhibit emergent unsafe behaviors not present in constituent models, yet exhaustive safety evaluation across all possible inputs remains intractable. Finally, the alignment properties of merged models may be more brittle than those of directly trained models, showing greater susceptibility to jailbreaking attacks (Djuhera et al., 2025).
@@
Safety-aware merging strategies have gained prominence in addressing adversarial vulnerabilities of merged models. Recent work on mitigating backdoor effects in multi-task model merging through safety-aware subspace identification (Djuhera et al., 2025) highlights growing recognition that naive merging may inadvertently propagate or amplify security vulnerabilities present in constituent models. Arora et al. (2024) show that model merging itself can serve as a defense mechanism, effectively sanitizing backdoored models through parameter averaging, a “free lunch” approach that eliminates backdoor triggers without requiring knowledge of the attack. Conversely, Merge Hijacking (Yuan et al., 2025) reveals that adversaries can craft malicious models specifically designed to inject backdoors through the merging process, highlighting the dual-use nature of merging for security. Merger-as-a-Stealer (Lu et al., 2025) further shows that model merging can be exploited to steal targeted personally identifiable information from aligned LLMs. SafeMERGE (Djuhera et al., 2025) proposes selective layer-wise model merging to preserve safety alignment in fine-tuned LLMs. Bias Vector (Shirafuji et al., 2025) introduces a systematic approach to mitigating biases in language models using the task arithmetic framework, showing that targeted bias task vectors can be constructed and negated to reduce multiple dimensions of social bias simultaneously. Personalized Soups (Jang et al., 2023) extends the alignment approach by enabling personalized LLM alignment via post-hoc parameter merging, allowing individual users to customize the alignment properties of merged models according to their specific value preferences. These approaches project merged parameters onto subspaces that preserve safety-critical behaviors while enabling capability transfer, often using activation analysis to identify safety-relevant parameter subspaces. WARM (Ram’e et al., 2024) shows the benefits of weight averaged reward models for improving alignment robustness. This represents a principled integration of security considerations into the merging pipeline.
The healthcare domain exemplifies the importance of alignment-preserving merging strategies, where domain-specialized models must maintain stringent safety properties including factual accuracy and appropriate uncertainty quantification. Developing uncertainty-aware merging protocols and factual consistency constraints that ensure clinical expertise transfer does not compromise safety guardrails remains an open challenge for medical model merging.
@@
Similarly, cybersecurity applications demand rigorous attention to the security implications of model merging, as merged models may exhibit unexpected vulnerabilities arising from the combination of individually benign model components. Adversarial robustness preservation during merging remains an underexplored direction. These domain-specific considerations become even more complex when merging must occur across organizationally or geographically distributed settings, where data privacy constraints and communication limitations introduce additional challenges that fundamentally reshape how model combination can be performed.
@@
The same arithmetic that enables safety injection also enables safety removal, since subtracting the alignment task vector $\tau_{\text{safe}}$ from a safety-tuned model can recover an unaligned variant, raising concerns about misuse (Ilharco et al., 2023). In addition, reward model merging via WARM (Ram’e et al., 2024) has shown that averaging independently trained reward models reduces reward hacking susceptibility, but the optimal number of reward models to merge and the conditions under which averaging preserves reward calibration remain theoretically uncharacterized. These dual-use considerations necessitate community-level governance frameworks for responsible model merging.